OpenStudy (anonymous):

please help, Indicate in standard form the equation of the line through the given points. K(6, 4), L(-6, 4)

OpenStudy (anonymous):

use this ... \[\frac{ y-y1 }{ x-x1 }=\frac{ y2-y1 }{ x2-x1 }\] plug in values and get result..after solving

OpenStudy (anonymous):

thanks